### v1

@@ 5 Application
Conversely, we want to know what are the observation points that have the most influence on the final estimated raked values. We look at the raked value for cause “injuries”, race group Black and New Castle County (denoted $\beta_{2,2,2}$). Then we compute the corresponding values of the gradient $\frac{\partial\beta_{2,2,2}}{\partial y_{i,j,k}}$ and $\frac{\partial\beta_{2,2,2}}{\partial s_{i}}$ for $i=0,\cdots,3$, $j=0,\cdots,5$ and $k=1,2,3$. The corresponding values of the gradient are shown in Figure 10. The most important observation point is the one with the same cause, race and county as the raked value considered. Increases in all-causes and all-races deaths will increase the raked value, while increases in other causes and races will decrease it. The initial values for the other counties will have a smaller influence on the raked value. For the margins, increases in all-causes, all-races state mortality and in all-races state mortality due to injuries will increase the raked value while increases in all-races state mortality due to other causes will decrease it.
@@
Finally, we compute the variance of the estimated raked values using two methods: First, we take $100$ samples of the detailed and aggregate observations, apply the raking procedure to each sample, and compute the sample mean and the sample variance of the corresponding raked values. Second, we compute the sample mean of the initial data set, apply the raking procedure to the sample mean, and use the delta method and the implicit function theorem to propagate the variance of the observations to get the variance of the raked values. In Figure 11, we looked at the raked values for Delaware, both sexes, and all age groups and compared the raked values and their standard deviation computed with both methods. We get very similar results with the delta method as we would have with a Monte Carlo technique that rakes all draws, gaining huge savings in compute.
Figure 11: Comparison between the means (left) and the standard deviations (right) of the raked values computed with the delta method or with all the samples. The method we developed gives nearly identical results as the Monte Carlo method that requires solving a raking problem for each draw of reported estimates and marginal totals.
@@ 6 Conclusions
In this paper, we have considered the multi-dimensional raking problem, where observations are adjusted to match marginals; for example, adjusting estimated deaths across race, county, and cause to match state all-race all-cause totals. We have provided asymptotically valid uncertainty estimates for inferred parameters that correctly account for uncertain constraints. To capture the uncertainty in both reported data and marginals, we leveraged the Lagrangian framework for optimization with linear constraints, and used the implicit function theorem to obtain the asymptotic uncertainty estimates using primal-dual optimality conditions. The approach obtains nearly the same uncertainty estimates as Monte Carlo estimation techniques that pass thousands of marginal ‘draws’ through the multi-dimensional raking estimation process, at the cost of a single solve.
@@ Acknowledgments
This work was funded by the Bill and Melinda Gates Foundation.

### v2

@@ 5 Application
Conversely, we want to know which observation points have the most influence on the final estimated raked values. We look at the raked value for cause “injuries”, race group Black and New Castle County (denoted $\beta_{2,2,2}$). Then we compute the corresponding values of the gradient $\frac{\partial\beta_{2,2,2}}{\partial y_{i,j,k}}$ and $\frac{\partial\beta_{2,2,2}}{\partial s_{i}}$ for $i=0,\cdots,3$, $j=0,\cdots,5$ and $k=1,2,3$. The corresponding values of the gradient are shown in Figure 11. The most important observation point is the one with the same cause, race and county as the raked value considered. Increases in all-causes and all-races deaths will increase the raked value, while increases in other causes and races will decrease it. The initial values for the other counties will have a smaller influence on the raked value. For the margins, increases in all-causes, all-races state mortality and in all-races state mortality due to injuries will increase the raked value while increases in all-races state mortality due to other causes will decrease it.
@@
Finally, we compute the variance of the estimated raked values using two methods: First, we take $100$ samples of the detailed and aggregate observations, apply the raking procedure to each sample, and compute the sample mean and the sample variance of the corresponding raked values. Second, we compute the sample mean of the initial data set, apply the raking procedure to the sample mean, and use the delta method and the implicit function theorem to propagate the variance of the observations to get the variance of the raked values. In Figure 12, we look at the raked values for Delaware, both sexes, and all age groups and compared the raked values and their standard deviation computed with both methods. We see evidence of sampling variation in the draws method, which makes sense since we only have 100 draws. The results obtained with the delta method are therefore both more accurate and far more efficient to compute compared to a Monte Carlo technique that rakes all draws.
Figure 12: Comparison between the means (left) and the standard deviations (right) of the raked values computed with the delta method or with all the samples. The new delta method gives more accurate results than the Monte Carlo method with a limited number of draws.
@@ 6 Conclusions
In this paper, we have considered the multi-dimensional raking problem, where observations are adjusted to match marginals; for example, adjusting estimated deaths across race, county, and cause in order to match state all-race all-cause totals. To capture the uncertainty in both estimated data and marginals constraints, we leveraged the Lagrangian framework for convex optimization with affine constraints, and used the implicit function theorem to obtain the asymptotically valid uncertainty estimates using primal-dual optimality conditions. The approach is orders of magnitude more computationally efficient than Monte Carlo techniques that pass realizations of observation draws and marginal draws through the multi-dimensional raking estimation process.
@@ Acknowledgments
This work was funded by the Bill and Melinda Gates Foundation. The Python scripts used to make the figures for the synthetic dataset can be found on the first author’s Github account (https://github.com/ADucellierIHME/JUQ_paper).